# Variational-Cartan Quantum Dynamics Simulations of Excitation Dynamics Linyun Wana, Jie Liub∗, Zhenyu Lia,b and Jinlong Yanga,b‡

### v1

@@ 3.3 Molecular Absorption Spectra
Designing organic electronic devices have attracted broad interest over past decades54. Acenes are an important ingredient in two classes of electronic devices: field-effect transistors (FETs, also knownas thin-film transistors, TFTs) and organic light-emitting diodes (OLEDs). The colour of acenes and Heteroacenes in the visible spectrum is primarily determined by the energy gap between their highest occupied molecular orbital (HOMO) and lowest unoccupied molecular orbital (LUMO). In principle, the active space that consists of the HOMO and LUMO is able to qualitatively describe the lowest electronic excitation of polyacenes. While, in order to quantatively describe their spectra, we need to include the dynamic correlation effect beyond the active space. As such, all the computed spectra of polyacenes have been shifted by -1.9 eV, which accounts for the difference between the complete active space (CAS) configuration interaction (CI) and CAS second-order perturbation theory 55.
Fig.  6: (a) Time evolution of the dipole moment for anthracene computed using the exact time-evoluation, VCQDS and VQDS approaches. (b) Absorption spectra simulated by the Fourier transformation of the dipole moments. The parameter $\gamma$ of the full width at half maximum is 0.001.
We apply the VCQDS algorithm to simulate molecular absorption spectra for a series of polyacenes, including naphthalene, anthracene, tetracene, and pentacene, which are polycyclic aromatic hydrocarbons composed of linearly fused benzene rings. The absorption spectra is an important tool to characterize the optical property of polyacenes. As the number of fused benzene rings increases, the first absorption peak exhibits an an obvious color change from blue to green to yellow due to the evolution of electronic structure 56.
@@
Figure. 7 shows the absorption spectra $\sigma_{\mathrm{abs}}(\omega)$ simulated using the VCQD approach for naphthalene, anthracene, tetracene, and pentacene. As the number of benzene rings increases, the wavelength of the signal gradually increases, which is consistent with the trend of redshift observed in the experimental spectral data 56. In addition, the intensity of the peaks gradually decreases from anphthalene to pentacene.
Fig.  7: Absorption spectra simulated using the VCQD approach for naphthalene, anthracene, tetracene, and pentacene. The parameter $\gamma$ of the full width at half maximum is 0.01. All the computed spectra have been shifted by -1.9 eV estimated for dynamic electron correlation.

### v2

@@ 3.2 Molecular Absorption Spectra
Designing organic electronic devices have attracted broad interest over past decades55. Acenes are an important ingredient in two classes of electronic devices: field-effect transistors (FETs, also knownas thin-film transistors, TFTs) and organic light-emitting diodes (OLEDs). The colour of acenes and Heteroacenes in the visible spectrum is primarily determined by the energy gap between their highest occupied molecular orbital (HOMO) and lowest unoccupied molecular orbital (LUMO). In principle, the active space that consists of the HOMO and LUMO is able to qualitatively describe the lowest electronic excitation of polyacenes. While, in order to quantatively describe their spectra, we need to include the dynamic correlation effect beyond the active space. As such, all the computed spectra of polyacenes have been shifted by -1.9 eV, which accounts for the difference between the complete active space (CAS) configuration interaction (CI) and CAS second-order perturbation theory 56.
Fig.  3: (a) Time evolution of the dipole moment for anthracene computed using the exact time-evoluation, VCQDS and VQDS approaches. (b) Absorption spectra simulated by the Fourier transformation of the dipole moments. The parameter $\gamma$ of the full width at half maximum is 0.001.
We apply the VCQDS algorithm to simulate molecular absorption spectra for a series of polyacenes, including naphthalene, anthracene, tetracene, and pentacene, which are polycyclic aromatic hydrocarbons composed of linearly fused benzene rings. The absorption spectra is an important tool to characterize the optical property of polyacenes. As the number of fused benzene rings increases, the first absorption peak exhibits an an obvious color change from blue to green to yellow due to the evolution of electronic structure 57.
@@
Figure. 4 shows the absorption spectra $\sigma_{\mathrm{abs}}(\omega)$ simulated using the VCQD approach for naphthalene, anthracene, tetracene, and pentacene. As the number of benzene rings increases, the wavelength of the signal gradually increases, which is consistent with the trend of redshift observed in the experimental spectral data 57. In addition, the intensity of the peaks gradually decreases from anphthalene to pentacene.
Fig.  4: Absorption spectra simulated using the VCQD approach for naphthalene, anthracene, tetracene, and pentacene. The parameter $\gamma$ of the full width at half maximum is 0.01. All the computed spectra have been shifted by -1.9 eV estimated for dynamic electron correlation.
3.3 Spin evolution of the Heisenberg model
We apply the hybrid Hamiltonian simulation approach to study one-dimensional Heisenberg models with competing interactions between nearest neighbor sites in periodic chains. The elementary excitation in the Heisenberg model is known as a magnon, which behaves as a boson after the Holstein-Primakoff (H-P) transformation. Simulating bosonic excitations is challenging using the diagonalization approach because there is no upper limit on the number of bosons. Alternatively, one can compute the dynamical correlation function to extract the spectra. The spin system is a natural candidate for simulation on a quantum computer 58, 59. Here, we require to compute spin susceptibilities, which can be defined as:
[…]
where $S_{z}^{i}(\omega)$ and $V_{j}(\omega)$ denote the Fourier transformations of the z-direction spin component on the $i^{th}$ qubit and applied electric fields in the z-direction on the $j^{th}$ qubit, respectively. In linear response theory, the susceptibility is closely related to a retarded Green function, so the correspoding relation for the two-point correlation functions ($C_{ij}(\omega)$) is 60:
$H=\sum_{\braket{i,j}}J_{x}X_{i}X_{j}+J_{y}Y_{i}Y_{j}+J_{z}Z_{i}Z_{j}.$ (35)
$e^{-iHt}=e^{-iJ_{x}tX_{1}X_{2}}e^{-iJ_{y}tY_{1}Y_{2}}e^{-iJ_{z}tZ_{1}Z_{2}},$ (36)
which differs the ground state of the antiferromagnetic Heisenberg model by an arbitrary global phase. We apply a $\delta$-kick external field on the first site, so the $\hat{D}$ is $Z_{1}$ and set field strength $E_{0}$ to $10^{-5}$. The time evolution of z-component of the spin on each site in the two-site Heisenberg model is shown in Fig. 6 (a). It is evident that the spin z-component exhibits a strict periodic oscillatory behavior between two sites and this oscillation satisfies convervation of spin z-component. Following Eq. 33 and Eq. 34, we can obtain the corresponding two-point correlation function between different sites, as shown in Fig. 6 (b) and (c), respectively. One can see that there is an evident symmetry and phase difference relationship between the two correlation functions. The phase difference between them accurately reflects the wave vector and the intensity of the spin wave, which can be obtained through Fourier transform to yield the corresponding magnon spectrum 58.
We also apply the hybrid Hamiltonian simulation approach to study the four-site antiferromagnetic periodic Heisenberg model. The corresponding geometric structure of the four-site periodic Heisenberg model is shown in Fig. 5 (b). The difference between two-site and four-site Heisenberg models lies in the non-commutativity between two certain terms in the four-site Heisenberg Hamiltonian. As discussed in Sec. 2.2, the Cartan decomposition can separate out the mutually commuting parts, which makes the time in the time evolution operator become a parameter in the circuit, and thus one can carry out the time evolution without Trotterization.
Fig.  5: Geometric structures of the Heisenberg model: (a) two-site (b) four-site. (c) Quantum circuit for implementing $\exp(-iHt)$. $H$ is the bare two-site Heisenberg model Hamiltonian. "$X"$ denotes Pauli ”$X$”, and $R_{\alpha}(\theta)$ denotes a rotation by $\theta$ on the $\alpha$-axis.
Fig.  6: (a) Time evolution of spins in the Z direction for the two-site antiferromagnetic Heisenberg model. Real and imaginary parts of (b) $\braket{S^{1}_{z}(t)S^{1}_{z}(0)}$ and (c) $\braket{S^{2}_{z}(t)S^{1}_{z}(0)}$ for the two-site antiferromagnetic Heisenberg model. $|J|$ is the coupling strength, $T$ is the simulation time.
Fig.  7: (a) Time evolution of spins in the Z direction for four-site antiferromagnetic Heisenberg model. (b) Magnon spectra of four-site antiferromagnetic Heisenberg model response to a $\delta$-kick external field. $|J|$ is the coupling strength, $q$ is the reciprocal lattice vector.